If 7 is added to a number the result is 30 times the reciprocal of the number. What is the number?
Can you please help me? Thanks in advane:)

Answer by nerdybill(7384)

If 7 is added to a number the result is 30 times the reciprocal of the number. What is the number?
Let x = the number
then
1/x = the reciprocal
.
From:"7 is added to a number the result is 30 times the reciprocal of the number" we get our equation:
7 + x = 30 * 1/x
7 + x = 30/x
multiplying both sides by x:
7x + x^2 = 30
x^2+7x = 30
x^2+7x-30 = 0
(x+10)(x-3) = 0
x = {-10,3}
The "number" can be either solution -- but, I would pick 3.